# CayleyPy RL: Pathfinding and Reinforcement Learning on Cayley Graphs

### v1

@@ 2.4. Novel method - combining diffusion distance with deep Q-learning
Part 1. Warm-up diffusion distance training. Use the diffusion distance approach for preliminary training the neural network. I.e. apply the steps 1,2 above: generate random walks to create training data, and train neural network to predict number of steps of random work.
Part 2. Modified DQN training. The second phase will relies on the Bellman equation strategy with mild modifications, but on somewhat different space exploration strategy comparing to conventional DQN - borrowed from the diffusion distance approach:
Step 1. Generate $N$ non-backtracking random walks trajectories all starting from a selected node $e$ (identity of a group). Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method.
Step 2. For the obtained nodes compute the Bellman equation predictions $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$ in the standard way.
Step 3. Clip the predictions by the number of steps of the random work since true distance is always smaller than number of steps. (Clip also predictions negative values).
Step 4. Run gradient descent for the neural network to minimize the loss function between clipped $t(g)$ (target) and neural network predictions.
Part 3. Graph search algorithm - beam search with neural network as an heuristics. After the neural network is trained we utilize the same approach to find a path by beam search and neural network as in our previous paper.
@@
The advantages of the proposed method are the following: there is no sparse reward problem - since we always start from the selected state, it always allows us to clip overestimated predictions, the warm-up phase allows to avoid long initialization process until when targets obtained by Bellman equation random since network starts from random weights - in our approach targets are always meaningful - first they are steps of random walks, at the second phase neural by itself produces meaningful targets since it was preliminary trained. We also suggest to use non backtracking random walks by rather clear reason - number of steps for non-backtracking is more similar to true distance. Conceptually that is related to the well-known phenomena - ”Non-backtracking random walks mix faster” [Alon07].
Benchmarks and analysis. Below we present the analysis of that proposal. We observe stable improvement of the method above comparing to diffusion distance approach, unfortunately the effect is not so big.
The first experiment is concerned not large groups, where one can compute by brute force the true graph distance and compare it with the predictions of the neural network. We see that for small sizes of the group methods learn practically ideally the distance on a graph, however for large groups correlation becomes not so perfect. Additional training by DQN always improves the preliminary warm-up training by diffusion distance procedure.
$n$ epochs: 5+0 epochs: 50+0 epochs: 5+50 epochs: 50+50
6 0.955 0.99 0.99 0.99
8 0.954 0.969 0.98 0.982
10 0.883 0.923 0.939 0.938
Table 1. n - permutation length. Spearman correlation of the true distance (obtained by brute force on million states) vs predictions of the neural networks trained by different methods. Conclusion Additional training with DQN improves results. Epochs A+B denotes: A - training epochs in warm-up diffusion distance part, B - epochs in DQN part.
The second experiments concerns larger group $S_{28}$ with LRX generators. We train the model and search for path by the beam search. The length of the path is the measure of performance. Experiment confirms that additional training by DQN improves the results. The combinatation of the 30+200 epochs training with 30-warmup and 200 - second stage (DQN) - is one the best which we observed for that case. Average length is 402, best is 382 which is not far from the expected ideal length 378. (In all experiments we searched path from the same element - conjecturally the longest element for LRX generators). For each experiment models for retrained from scratch, thus due to random nature of the random walks models are significantly different.
Experiment epochs: epochs: epochs: epochs:
30+200 30+0 100+0 0+230
1 490 None 446 None
2 396 658 516 None
3 382 None None None
4 408 650 436 None
5 428 706 1046 None
6 396 None None 508
7 426 890 614 None
8 None 598 None None
9 392 464 524 None
10 None 626 452 434
Median 402 650 516 455.5
Table 2. Lengths of solutions obtained by different approaches. $S_{28}$ LRX generators. 10 experiments to decompose the ”longest” element with expected ideal length 378. Conclusion: additional training by DQN improves the results - both in length and probability to find a path. Pure DQN training is not performing well.
The analysis can be found in notebook.
3. Benchmarks
Goal - $S_{110}$. DeepMind’s success with Go game with $10^{170}$ states - creates a challenge to achieve similar results in the other fields in particular for Cayley graphs, at the moment we are somewhat far from that, but there are plenty ideas not yet incorporated in our CayleyPy project and optimistically such a goal is achievable in near horizons. Moreover in the present study we achieve pathfinding for such scale groups, but unfortunately not by fully zero prior knowledge. Our zero prior knowledge approaches achieves pathfiding for LRX generators of $S_{n}$ a bit above $30$, but at the moment $40$ is a not achieved despite bunch of neural networks we tried. What is surprising that almost all approaches we tried stuck slightly above $n=30$ all in the same region, with small variations. It is significantly better than results of computer algebra system GAP which achieves only about $n=20$, which much more lengths and timings.
3.1. One line of code to get orders of magnitude improvement from: $S_{30}$ to $S_{100}$.
if action == ’X’ and current-state[0] < current-state[1]: continue The condition used in the beam search part. Its logic is rather clear - if the contents of positions 0 and 1 are already sorted - it is logical not to swap them i.e. not to apply $X$ generator.
Surprisingly such a simple conditions is a game changer. Inserting it and using it with exactly the same pipelines which performed for around $n=30+$ we are able to find paths for $n=100$ and (apparently more - just stopped by 16G RAM limit). So the size of graphs changes from $10^{3}0+$ to $10^{1}50+$. It works the same for all models: perceptrons, gradient boostings, convolutional neural networks.
The challenge is how to avoid such prior knowledge method and extend it to other generators.
It should also be noted that this condition prevents to find optimal paths. And for small sizes of $n\leq 15$ it typically leads to a bit longer paths, but for higher values of $n$ it is a relief.
Examples of lengths found by that method are presented at table 3 below:
$S_{n}$ 100 90 80 70 60
length 16632 33459 7510 8115 3650
ideal length 4950 4005 3160 2415 1770
Time 8h 43m 3h 57m 1h 5m 52m 3h
Nb version 255 242 241 233 -
Table 3. n - permutation length. Lengths of solutions for different $n$ obtained by ML model combined followed by beam search with ”X-condition”. (Not completely zero prior knowledge approach).
Computation for $n=100$ took 8h 43m 53s GPU P100. Beam width ${2}^{20}$, 30 epochs of training. ( Notebook version 255), for other cases we used beam size not more than ${2}^{18}$, that is why for $n=90$ length is much bigger. The data for $n=60$ is obtained by the gradient boosting model, with slightly better lengths, but more timing, in the other cases single layer MLP model used. In all cases we decompose the conjecturally longest element of expected ideal length $n(n-1)/2$.
3.2. ”Overcome the GAP” with zero prior knowledge AI.
@@ 3. Benchmarks
We use LRX Cayley graph of $S_{n}$ and its conjecturally longest element of expected length $n(n-1)/2$ as a benchmark. Below we demonstrate that AI methods significantly outperform classical methods of computer algebra system GAP.
$n$ GAP Ideal AI
Length Length n(n-1)/2 Length
@@
Table 4. n - permutation length. Comparison of GAP vs AI methods (zero prior knowledge). On the conjecturally longest element of LRX Cayley graph. (Its expected ”ideal” length is $n(n-1)/2$.)
@@
For $n=20$ GAP timing is 41min 18s, while AI methods can find results much faster. For example, simple neural network: notebook version 423: 3m 48s GPU P100. GAP benchmarks has been performed in the notebook.
$n$ Ideal AI
Length n(n-1)/2 Length
22 231 231
23 253 253
24 276 276
25 300 300
26 325 325
27 351 351
28 378 378
29 406 408
30 435 437
31 465 478
32 496 728
33 528 642
Table 5. n - permutation length. Comparison of GAP vs AI methods.
The AI results above are aggregated results from experiment with different models and approaches. In particular based on gradient boostings and neural networks, all of them quite stably and perfectly solve cases with $n\leq 30$, but going beyond that requires efforts for all our approaches.It is surprising that all types of models have the problem at the same range of $n$ and that length is not becoming gradually longer and longer, but it stays almost perfect, and then quite sharply pipeline loses ability to find a path at all. For some cases we use several reruns - if the path not found via single run, it sometimes be found by multiple relaunches of the pipeline. For example $n=32$ has been solved with our modified DQN strategy with 30 warmup epochs and 1000 DQN epochs (notebook), time: 2h 53m 53s GPU P100.
@@
The tests of different architectures and parameters is discussed below. The main conclusion that we observe that rather simple models like MLP and gradient boostings currently perform better, than more advanced: transformers, convolutional neural networks etc. One exception is one of our transformer models that can find paths up to $n=100$, however it is not zero prior knowledge - it uses features specific to LRX generators. Dependence on various parameters was explored in our previous paper  [Chervov25] for the case of the Rubik’s cube group, the conclusions were that beam size is one the most important parameter the length of solution is almost linearly improving on logarithm of beam size, deeper neural networks are better than wider especially, but the effect of depth is better seen on small for higher cubes, we also observed stagnation of performance with increase of training set.
@@
For the LRX Cayley graph these phenomenons are less strongly expressed, or even negligible. Most experiments with deep neural networks have not lead to significant improvements over simple one layer perceptron with 128 neurons. Beam size effect is present, in but less expressed (see table below).
Nonbacktracking random walks have quite strong effect. In particular one can train models in just single epoch within a minute for e.g. $S_{20}$. We actually use improved version of the nonbacktracking based on the following ideas: 1) we forbid not only 1 step in history but typically 32 steps (it is a parameter) 2) we consider a bunch of trajectories and ban states for all of them at once 3) we ban not only visited states but their neighbors. The price for that is slowing down the generation, but that is not so important since gradient descent takes much longer time, and overall there is no significant slow down.
Beam Solution Length Solution Length
Width Nonbacktracking RW Simple RW
2**10 None None
2**12 252 None
2**14 202 None
2**16 194 None
2**18 194 None
2**20 190 (Ideal) None
Table 6. Without non-backtracking system cannot perform well on single epoch training. Effect of beam size is clear, but not so strong.
We generate 10000 trajectories at once each of the length 190, by the generalized nonbacktracking procedure above and train the neural network with batch size 1024 just for single epoch. Entire time for each computation - both training and beam search are 1-5 minutes, experiments can be found in versions 427-430 of the notebook.
Theoretical reason why non-backtracking is helpful is clear. Imagine our graph is a tree, then even for non-backtracking random walk number of steps coincide with the true distance (length of the shortest path). For complicated graphs it is not so, but still it more close to true distance and has less variability. I.e. the situation when the same node can be achieved in different number of steps - the fact which equivalent to having a noise in the training data and which makes models less precise.
@@
CNN. We also analyse convolutional neural networks - however they have slower training process. They show similar results comparing to perceptrons e.g. the length of the solution for $n=16$ is 122 (ideal is 120), for $n=20$ is 216 (ideal is 190). The basic version of the architecture uses two 1D CNN layers with Batch Normalization and Dropout to avoid overfitting. One of the successful variations is using temporal gated CNN. The first part of the architecture is the residual block - several convolutional blocks with gating based on sigmoid activation function. The second part of the architecture is stacking several layers of residual blocks with gradually increased dilation to increase the size of the receptive field. Notebooks: 1, 2, 3,4.
@@
Transformer. We also peformed experiments with a transformer model, however it was able to solve only not so big groups like $n\leq 15$. That model incorporated standard Transformer components, including multi-head self-attention, positional encoding, feed-forward layers, and dropout to mitigate overfitting. However, the model struggled to converge to a satisfactory solution, even for relatively short sequences. This issue became increasingly severe as the sequence length grew. For instance, with size = 9 (corresponding to path length = 107 and ideal = 36) or n-size = 13 (path length = 332, ideal = 78), increasing the projection dimension to 512 or 1024, adding more layers, or training for more epochs failed to address the underlying challenges. Notebooks: 1, 2.
@@
Transformer based on LRX specific features. The features inspired by Hamming distance were used: difference between an element and its position; difference with the left neighbor; difference with the right neighbor. A standard multihead attention with one encoder layer and mlp was used to estimate the number of steps required to reach the ordered state. The transformer demonstrated the ability to handle sequences up to N=15 without using beam search. For larger N, beam search is employed, and the maximum N for which results were obtained was 100. The model and features only indirectly depend on N, therefore, when trained on micro batches constructed as random walks from an ordered permutation for N*(N-1)/2 moves, it can be used for different N values, both larger and smaller. (Notebook).
@@
Gradient boostings. We also applied CatBoost and XGboost in our diffusion distance pipeline, surprisingly they perform not worse than neural networks (it was not the cases for Rubik’s cube - where boosting were not able to solve 4x4x4, 5x5x5 cubes), and even giving a bit shorter solutions. (Code).
3.4. Reinforcement learning on ”small graphs”
At the preliminary stage of the research we performed experiments with RL/dynamical programming methods for small graphs. In particular the question was how the initialization influence the performance. That is important to understand in view of our proposal to use initialization coming from the diffusion distance. These experiments revealed interesting phenomenons presented at figure 2. For some specific initialization the performance first degrades, and only after that achieves the convergence to the true distance which is guaranteed by the general theory. The number of steps for convergence is not more than diameter of the graph is initialization is positive, however can be larger in presence of negative values. Initialization coming from the machine learning model (”catboost”) trained on diffusion distance (same as ”warmup step” in our proposal) improves the convergence, though not in a radical manner.
Figure 2. Pearson correlation between $d_{i}(s)$ and $d(s)$ on each iteration of the DP algorithm, $\epsilon=10^{-3}$

### v2

@@ 2.4. Combining diffusion distance with deep Q-learning
Part 1. Warm-up diffusion distance training. Apply diffusion distance approach for pre-training the neural network, i.e. steps (1) and (2) above: generate training data via random walks, and train a neural network $f_{\boldsymbol{\theta}}(s)$ to predict the number of steps of the random walk.
Part 2. Modified DQN training. We use the deep Q-learning algorithm 2 with two enhancements:
  • the random subset of nodes consist of all states visited by $N$ random walks of size $K_{\max}$;
  • the result of the Bellman update $1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$ is clipped to the interval $[0,k]$ where $k$ is the time step when the state $s$ was visited for the first time.
The pseudo code for this modified version of DQN is presented in Algorithm 4.
Part 3. Guided beam search. Here we apply step (3) from the diffusion distance approach described above (Algorithm 3).
Algorithm 4 Modified deep Q-learning
1:Hyperparameters: $M$ (number of epochs), NN (number of random walks), $K_{\max}$ (length of random walks), $f_{\boldsymbol{\theta}}(s)$ (distance estimator)
2:for $i=0$ to $M$ do
3:  Generate $N$ random walks of size $K_{\max}$ from the target node $e$
4:  Collect dataset $\mathcal{D}=\{(s,k)\}$ of states $s$ visited by random walks at time $k$
5:  for each $(s,k)\in\mathcal{D}$ do
6:   $d(s)\leftarrow 1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$ (update based on Bellman equation (2))
7:   $d(s)\leftarrow\min\{k,\max\{0,d(s)\}\}$ (clipping)
8:  end for
9:  Calculate MSE loss $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{|\mathcal{D}|}\sum\limits_{\mathcal{D}}\big(d(s)-f_{\boldsymbol{\theta}}(s)\big)^{2}$
10:  Calculate $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$ and do the gradient step
11:end for
@@
There are several advantages of the proposed method. First, it avoids the sparse reward problem: since we always start from the selected state, we can reliably clip overestimated predictions. Second, the warm-up phase helps avoid the long initialization typical for standard DQN, where targets from the Bellman equation are noisy because the network starts with random weights. In our approach, the targets are meaningful from the beginning: in the first phase, they come from random walk steps; in the second phase, the network itself produces useful targets, as it has already been pre-trained. Finally, we use non-backtracking random walks for a clear reason: the number of steps in such walks better reflects the true distance. Conceptually, this relates to the well-known phenomenon ”Non-backtracking random walks mix faster” [Alon07].
3. Analysis and results
DeepMind’s success in mastering the game of Go, which has approximately $10^{170}$ possible states, poses an inspiring challenge: achieving similar breakthroughs in other domains, such as Cayley graphs. At present, we are still a distance from this goal. However, many promising ideas have yet to be implemented in our CayleyPy project, and with continued development, reaching such a milestone may be possible in the near future.
In this study, we demonstrate successful pathfinding in very large LRX Cayley graphs of $S_{n}$. We use its conjecturally longest element (1) of expected length n(n1)/2n(n-1)/2 as a benchmark. Our zero prior knowledge approach can find the paths from $\ell_{n}$ to $e$ for $n\leqslant 44$. This is a significant improvement over the performance of the GAP computer algebra system, which can only handle up to around $n=20$, and with much longer path lengths and runtimes. Introducing some domain knowledge (see Section 3.2 below) drastically improves the performance to $n=100$ or even more.
3.1. ”Overcome the GAP” with zero prior knowledge AI
@@ 3. Analysis and results
Here we demonstrate that the diffusion distance method significantly outperform classical methods of computer algebra system GAP.
$n$ GAP length $\frac{n(n-1)}{2}$ DD length
@@
Table 2. Comparison of GAP and diffusion distance ($D$) method for the conjecturally longest element of LRX Cayley graph with expected length n(n1)/2n(n-1)/2.
@@
For $n=20$, the GAP timing is 41min 18s, while our methods can find results much faster. For example, the diffusion distance method with simple neural network requires only 3m 48s of calculations on GPU P100 (notebook version 423). GAP benchmarks have been performed here.
$n$ $\frac{n(n-1)}{2}$ DD length $n$ $\frac{n(n-1)}{2}$ DD length
22 231 231 31 465 478
23 253 253 32 496 508
24 276 276 33 528 642
25 300 300 34 561 667
26 325 325 36 630 834
27 351 351 38 703 759
28 378 378 40 780 926
29 406 408 42 861 909
30 435 437 44 946 1060
Table 3. Lengths of solutions produced by diffusion distance ($D$) method for the conjecturally longest element of LRX Cayley graph with expected length n(n1)/2n(n-1)/2.
The DD length columns in Tables 2 and 3 are aggregated results from experiments with different guiding models based on gradient boostings and neural networks. All of them solve cases with $n\leq 28$ quite stably; moreover, after some hyperparameter tuning, they find the shortest paths. Going beyond the limit $n=28$ requires some effort. For $n\leq 44$ a path can be occasionally found by multiple relaunches of the pipeline, however, its length is no longer optimal.
3.2. X-trick.
Surprisingly, adding one line of code allows us to drastically increase the size of solvable graph, from $S_{40}$ to $S_{100}$. Namely, add the following condition in line 9 of the beam search algorithm 3: if $s_{0}<s_{1}$ then do not include $X\circ s$ in $N(s)$. The logic is straightforward: if the first two elements of the permutation are already sorted, it makes little sense to swap them by applying the generator X=(0 1)X=(0\ 1).
This X-trick turns out to be a game changer. By adding it to the pipelines that succeeded for $n$ about $40$, we were able to find paths for $n=100+$ (and potentially even larger values, as we were limited by the 16GB RAM limit). This represents a shift from working with Cayley graphs of size $10^{30+}$ to $10^{150+}$. This trick worked universally across model types, including perceptrons, gradient boosting machines, and convolutional neural networks. The key challenge now is how to eliminate such prior knowledge and generalize the method to other types of generators. However, X-trick may prevent the discovery of optimal paths, especially for small values of $n\leq 15$.
Examples of lengths found by the diffusion distance method with X-trick are presented in Table 4 below:
$n$ 100 90 80 70 60
Solution length 16632 33459 7510 8115 3650
Diameter (proposed) 4950 4005 3160 2415 1770
Time 8h 43m 3h 57m 1h 5m 52m 3h
Notebook version 255 242 241 233 -
Table 4. Lengths of solutions for different $n$ obtained by ML model followed by beam search with X-trick. (Not completely zero prior knowledge approach).
Computation for $n=100$ took 8h 43m 53s GPU P100. Beam width ${2}^{20}$, 30 epochs of training ( Notebook version 255). For other cases we used beam size $\leq{2}^{18}$, that is why for $n=90$ length is much larger. The data for $n=60$ is obtained by the gradient boosting model, with slightly better lengths, but more timing, in the other cases single layer MLP model used. In all cases we decompose the conjecturally longest element of expected length n(n1)/2n(n-1)/2.
3.3. Diffusion distance + modified DQN
We observe that replacing the training stage with the DQN algorithm 4 gives a consistent improvement over the diffusion distance approach; however, the effect is not very large. Below we provide two illustrative experiments.
$n$ epochs: $5+0$ epochs: $50+0$ epochs: $5+50$ epochs: $50+50$
6 0.955 0.99 0.99 0.99
8 0.954 0.969 0.98 0.982
10 0.883 0.923 0.939 0.938
Table 5. Spearman correlation between the true distances and the predictions of the neural network trained for AA epochs during the warm-up diffusion distance phase and $B$ epochs during the DQN phase (epochs: $A+B$). Additional training with DQN leads to improved results.
The first experiment (see Table 5) focuses on relatively small groups, where the true graph distance can be computed by brute force and compared with the neural network’s predictions. For small group sizes, the method learns the graph distances almost perfectly. However, for larger groups, the correlation between predictions and true distances becomes less accurate. Still, additional training using the modified DQN always improves upon the initial warm-up phase based on the diffusion distance procedure.
Warm-up epochs DQN epochs success rate shortest path median path
30 200 80% 382 402
30 0 70% 464 650
100 0 70% 436 516
0 230 20% 434 471
Table 6. Path lengths for decomposing the longest element in the LRX Cayley graph of $S_{28}$. Results in each row are based on 10 independent runs. Additional training with DQN improves both the average path length and the success rate of finding a path. Pure DQN training alone performs poorly.
The second experiment concerns the larger group $S_{28}$ with LRX generators. We train the model and use beam search to find a path, where the path length serves as the performance metric. The results provided in Table 8 confirm that additional training with DQN improves performance. The combination of $30$ warm-up and $200$ DQN epochs produces one of the best outcomes observed in this setting. The median (shortest) path length $402$ ($382$) is very close to the expected ideal length of $378$. For each run, models were retrained from scratch; thus, due to the randomness of the walks, the resulting models differ significantly. The analysis can be found in the Kaggle notebook.
3.4. Non-backtracking
We exploit non-backtracking idea in two ways - first we can use non-backtracking random walks, and the second we can use non-backtracking in beam search. The second way shows quite a good effect.
history size, $m$ success rate median length min length
2 $40.41\%$ $476$ $380$
4 $40.10\%$ $474$ $382$
8 $40.32\%$ $471$ $386$
16 $37.40\%$ $470$ $382$
32 $37.47\%$ $470$ $378$
40 $36.16\%$ $480$ $390$
48 $37.39\%$ $465$ $380$
128 $36.45\%$ $476$ $382$
254 $37.85\%$ $468$ $382$
400 $37.40\%$ $474$ $384$
512 $36.58\%$ $476$ $384$
Table 7. Performance of the $D$ method with non-backtracking random walks ($n=28$).
Beam Solution Length
Width Non-backtracking RW
$2^{10}$ None
$2^{12}$ 252
$2^{14}$ 202
$2^{16}$ 194
$2^{18}$ 194
$2^{20}$ 190 (Ideal)
Table 8. Influence of beam width and non-backtracking random walks for solving the longest element of the LRX Cayley graph $S_{20}$.
The theoretical reason why non-backtracking is helpful is clear for the random walks. Imagine our graph is a tree, then for a non-backtracking random walk, the number of steps coincides with the true distance (length of the shortest path). Thus diffusion distance coincides with the standard distance. For complicated graphs, this is not the case, but still, it forces the diffusion distance to be closer to true distance and to have less variability. Here by variability we mean the situation when the same node can be achieved in a different number of steps. This is equivalent to having noise in the training data which makes models less precise. Similar for beam search non-backtracking mitigates stagnation in local minimums.
history size $m$ success rate median length min length
2 $17.65\%$ $430$ $384$
4 $17.08\%$ $432$ $390$
8 $21.37\%$ $440$ $382$
16 $30.07\%$ $446$ $388$
32 $38.55\%$ $473$ $384$
64 $43.29\%$ $495$ $380$
128 $56.95\%$ $517$ $384$
256 $60.61\%$ $588$ $384$
512 $67.33\%$ $582$ $384$
1024 $9.71\% 66$3663 $382$
Table 9. Performance of the $D$ method with non-backtracking beam search ($n=28$). The path is found more often as the history size grows.
Injecting non-backtracking into the beam search stage (Algorithm 3) allows to increase the success rate: the greater is the number of banned states not included in the beam, the higher is the probability of finding a solution (see Table 9). On the other hand, the median solution length tends to increase when the history size is large. But that cannot be interpreted as performance degradation, rather it means that pipeline can solve more difficult cases, and hence getting large lengths.
@@ 3.4. Non-backtracking
We observe that rather simple models like MLP and gradient boostings currently perform better than more advanced ones: transformers, CNNs, etc. An exception is one of our transformer models that can find paths up to $n=100$, however, it is not a zero prior knowledge model since it uses features specific to LRX generators.
@@
The dependence on various parameters was explored and discussed in our previous paper  [Chervov25] for the case of the Rubik’s cube group. The beam size proved to be one of the most important parameters: the length of solution is almost linearly improving on logarithm of beam size. Another observation is that it is more important for a neural network to be deep than wide, however, the effect of depth is better seen for large groups. We also noted the performance stagnation as size of the training set increased.
@@
For the LRX Cayley graph these phenomenons are less strongly expressed, or even negligible. Most experiments with deep neural networks have not lead to significant improvements over simple one layer perceptron with 128 neurons. In particular, one can train models in just a single epoch within a minute, e.g., for $S_{20}$. The beam size effect is still present but not as pronounced (see table 8).
@@
CNN. Convolutional neural networks have demonstrated a slower training process. Their results are similar to those of perceptrons, e.g., length of the solution for $n=16$ is 122 (ideal is 120), for $n=20$ is 216 (ideal is 190). The basic version of the architecture uses two 1-d CNN layers with batch normalization and dropout to prevent overfitting. One of the successful variations is using temporal gated CNN. The first part of the architecture is the residual block - several convolutional blocks with gating based on sigmoid activation function. The second part of the architecture is stacking several layers of residual blocks with gradually increased dilation to increase the size of the receptive field. Notebooks: 1, 2, 3,4.
@@
Transformer. We also performed experiments with a transformer model. However, it was able to solve only not-so-large groups like $n\leq 15$. That model incorporated standard transformer components, including multi-head self-attention, positional encoding, feed-forward layers, and dropout to mitigate overfitting. However, the model struggled to converge to a satisfactory solution, even for relatively short sequences. This issue became increasingly severe as the sequence length grew. For instance, with $n=9$ (corresponding to path length = 107 and ideal = 36) or $n=13$ (path length = 332, ideal = 78), increasing the projection dimension to 512 or 1024, adding more layers, or training for more epochs failed to address the underlying challenges. Notebooks: 1, 2.
@@
Transformer based on LRX specific features. The features inspired by Hamming distance were used: difference between an element and its position; difference with the left neighbor; difference with the right neighbor. A standard multihead attention with one encoder layer and MLP was used to estimate the number of steps required to reach the target state. The transformer demonstrated the ability to handle sequences up to $n=15$ without using beam search. For larger $n$, beam search is employed, and the maximum $n$ for which results were obtained is 100. The model and features only indirectly depend on $n$, therefore, when trained on micro batches constructed as random walks from the identity permutation for n(n1)/2n(n-1)/2 moves, it can be used for different values of $n$, both larger and smaller. (Notebook).
@@
Gradient boosting. We also applied CatBoost and XGboost in our diffusion distance pipeline. Surprisingly, they perform not worse than neural networks (it was not the cases for Rubik’s cube where boostings were not able to solve $4\times 4\times 4$, $5\times 5\times 5$ cubes), and even produce a bit shorter solutions. (Code).
More simulation results for various parameters can be found in spreadsheet here, or on google sheets.